OpenStudy (anonymous):

2√3-1 3√3-2 a.) -9√3 b.) 16+1√3 c.) 13√3 d.) 20-7√3

sam (.sam.):

You mean this ? \[2\sqrt{3}-13\sqrt{3}-2\]

sam (.sam.):

2~(3)-13~(3)-2 According \to the distributive property, for any numbers a, b, and c, a(b+c)=ab+ac and (b+c)a=ba+ca. \[\text{Here,} \sqrt{3} ~~\text{is a factor of both}~~ 2\sqrt{3}~~ and -13\sqrt{3}.\] \[(2-13)\sqrt{3}-2\] to add integers with different signs, subtract their absolute values and give the result the same sign as the integer with the greater absolute value So, subtract the absolute values of 2 and -13 and give the result the same sign as the integer with the greater absolute value. \[-11\sqrt{3}-2 \]

OpenStudy (anonymous):

(2√3-1)(3√3-2)

sam (.sam.):

(2~(3)-1)(3~(3)-2) FOIL 6\sqrt(3)^(2)-7\sqrt(3)+2 18-7sqrt(3)+2 20-7sqrt(3)

OpenStudy (anonymous):

Thank you!